OpenStudy (anonymous):

The table shows a direct variation. What is the constant of variation? x y 4 14 4.3 15.05 5 17.5

OpenStudy (anonymous):

-_-

hartnn (hartnn):

for direct variations, we have \(y = kx\) where k is the constant of variation. so, k = y/x

hartnn (hartnn):

just take any pair say x = 4, y =14 and get k = y/x = .. ?

OpenStudy (anonymous):

i dont get it

hartnn (hartnn):

which part did u not get?

OpenStudy (anonymous):

y=14

hartnn (hartnn):

y = 14 when x= 4 so k = y/x = 14/4 = ... ?

OpenStudy (anonymous):

y=14

OpenStudy (anonymous):

is that it

OpenStudy (anonymous):

i am in a rush

hartnn (hartnn):

what did u get for k = 14/4 = ..?

OpenStudy (anonymous):

4=...... i dont know

hartnn (hartnn):

hint : 14 = 2*7 4 = 2*2 so 14/4 = 2*7/2*2 what gets cancelled out?

OpenStudy (anonymous):

i am realy in a rush i need to finish in 10 mins

hartnn (hartnn):

why 10 mins?

OpenStudy (anonymous):

tesing

hartnn (hartnn):